OpenStudy (anonymous):

how to solve 2+cos^2(2 theta) =3 sin^2(2 theta)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2+\cos^2(2θ)=3\sin^2(2θ)}\) \(\normalsize\color{blue}{3-1+\cos^2(2θ)=3\sin^2(2θ)}\) \(\normalsize\color{blue}{3-(1-\cos^2(2θ)~~~~~~)=3\sin^2(2θ)}\) \(\normalsize\color{blue}{3-(\sin^2(2θ)~~~~~~)=3\sin^2(2θ)}\) \(\normalsize\color{blue}{3-\sin^2(2θ)=3\sin^2(2θ)}\) \(\normalsize\color{blue}{3=3\sin^2(2θ)-\sin^2(2θ)}\) \(\normalsize\color{blue}{3=2\sin^2(2θ)}\) this is how I am thinking to start looks stupid though...

OpenStudy (anonymous):

It says there should be 8 solutions

OpenStudy (phi):

you could use 1 - cos^2 = sin^2 to replace the sin^2 with a cos^2 then collect terms.

OpenStudy (anonymous):

Sorry I'm not really good at trig. Can u show the steps please

OpenStudy (solomonzelman):

phi, I see using `sin²x+cos²x=1` it the other way...

OpenStudy (phi):

\[ 2+\cos^2(2 \theta) =3 \sin^2(2 \theta) \] replace the sin^2(2θ) with (1-cos^2(2θ) ) can you do that ?

OpenStudy (phi):

yes, Solomon, there are a few ways to simplify this.

OpenStudy (solomonzelman):

yes, I thought of using that sin(x+x) and cos(x+x)

OpenStudy (phi):

replace means "erase" sin^2(2θ) and put in the new expression (1 - cos^2(2θ))

OpenStudy (anonymous):

So now it's 2+cos^2(2theta) = 3(1-cos^2(2theta)) now what

OpenStudy (phi):

distribute the 3 on the right hand side

OpenStudy (anonymous):

3-3cos^2(2theta)=2+cos^2(2theta)

OpenStudy (phi):

now we want the cos^2 on the same side. Add +3cos^2(2theta) to both sides and while you are at it, add -2 to both sides

OpenStudy (anonymous):

I simplified and got 0=4cos^2(2theta) then what would I do next

OpenStudy (phi):

close, but you should not get 0 on the left.

OpenStudy (phi):

if you add -2 to both sides, the left side is 3 - 2 (which is not 0)

OpenStudy (anonymous):

1=4cos^2(2theta) ?

OpenStudy (phi):

ok, now divide both sides by 4

OpenStudy (anonymous):

I simplified and got 1/2 = cos 2 theta

OpenStudy (phi):

yes, but it's ± ½ = cos(2 theta)

OpenStudy (phi):

I can see how to get 4 solutions out of that but it will take some thought to find the other 4

OpenStudy (anonymous):

Ok I know how to solve from here. Thank you so much, I've been trying to solve this problem for days!

OpenStudy (phi):

so far we have cos(2x)= +½ cos(2x) = -½ we can use cos(2x)= cos(x+x) = cos(x)cos(x) - sin(x)sin(x) or cos(2x) = cos^2(x) - sin^2(x) or using sin^2 = (1-cos^2(x)) we get cos(2x) = 2 cos^2(x) -1 use that in the two equations

OpenStudy (anonymous):

I got theta = pi/6, 5pi/6, pi/3, 2pi/3, 7pi/6, 11pi/6, 8pi/6, and 10pi/6 is that correct ?

OpenStudy (phi):

probably! try them in the original equation.

OpenStudy (phi):

yes, they all work